OpenStudy (anonymous):

An ice cube that measures 2.5 centimeters on each side contains how many cubic centimeters of ice? Include correct units with your solution.

OpenStudy (anonymous):

Volume of a cube is given by V = s^3 where s is the length of the side. Thus, V = (2.5)^3 = 15.625 Therefore, the ice cube contains 15.625 cubic centimeters of ice.

OpenStudy (anonymous):

That is correct.

OpenStudy (anonymous):

thanks